OpenStudy (anonymous):

I need help with fractions. 2/5+3/10

OpenStudy (anonymous):

Can you find a common denominator?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Common denominator would be the same for both fractions... can you express 2/5 as some number of 10ths, like A/10? In other words, 2/5 = A/10 what should A be?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

right, so you can re-write it all like this: 2/5 + 3/10 = 4/10 + 3/10 = 7/10

OpenStudy (anonymous):

Okay cool. Can you help me with another?

OpenStudy (anonymous):

Gotta run in a sec, but post it as a new question and I will give it a try

OpenStudy (anonymous):

Find x x^2(3x+2)(2x-5)^3

OpenStudy (anonymous):

Whoa :) Not fractions anymore...

OpenStudy (anonymous):

Oh sorry x^2(3x+2)(1/2x-5)^3

OpenStudy (aripotta):

lol

OpenStudy (anonymous):

Expand out 1/(2x-5)^3 as 1 [ (2x-5)(2x-5)(2x-5) } unless you know the shortcut rule for expanding a "cube" expression

OpenStudy (aripotta):

is it 1/(2x-5)^3 or (1/2x-5)^3?

OpenStudy (anonymous):

@AriPotta can you take this? I actually have to step away for a bit... I could have done one more common denominator fraction problem, but I can't stick around for this... @SarahQT345 , I will be back on later, sorry to run out mid-problem... good luck...

OpenStudy (aripotta):

or..does it not matter..?

OpenStudy (aripotta):

eh..i honestly don't really know what to do lol

OpenStudy (aripotta):

:l i feel like i just got left with like a dozen kids all under the age of 4.

OpenStudy (aripotta):

um...ok. x^2(3x+2)(1/2x-5)^3 i would do the distributive property. distribute the x^2 into the parentheses

OpenStudy (aripotta):

do you know how to distribute?

OpenStudy (anonymous):

Im in 5th grade. My brother gave me his homework and said that if I didn't do it that he'd beat me up.

OpenStudy (aripotta):

oh. really? that's not nice. what grade is he in?

OpenStudy (aripotta):

i can beat him up for you ;)

OpenStudy (anonymous):

12th. He has a gun.

OpenStudy (aripotta):

.......................oh.

OpenStudy (anonymous):

Please. I need help.

OpenStudy (aripotta):

i don't know how to do it :c

OpenStudy (aripotta):

i've been trying, i just don't understand what to do :/

OpenStudy (anonymous):

Well, I hope I survive through the week. You see, next week, it's all on Emily's shoulders. Emily is his girlfriend. She has been in Italy for the past week and this one. She normally does his homework with the same standards. If it doesn't get done, bad things happen. It kinda reminds me of Daddy when he drinks to much.

OpenStudy (aripotta):

i'm really sorry :/

OpenStudy (anonymous):

I think Daddy is home from the bar.

OpenStudy (anonymous):

x^2(3x+2)(2x-5)^3 what is this equal to? is this equal to 0?

OpenStudy (anonymous):

I don't have much time. I think Daddy is coming up stairs.

OpenStudy (anonymous):

x will have zeros at 0 and around -2/3 and 10, assuming your equation is set equal to 0

OpenStudy (anonymous):

Normally when Daddy comes home this late he tries to touch me.

OpenStudy (aripotta):

...

OpenStudy (anonymous):

you are revealing too much of your private life. You need professional help.

OpenStudy (anonymous):

My dad says he is a professional malastulator.

OpenStudy (anonymous):

...Like i said. professional help.

OpenStudy (aripotta):

if you're concerned about him touching you again, you need to find some time to call 9-1-1...like when you're brother's at school and you're dad's at the bar. or even talk to one of your teachers about it.

OpenStudy (aripotta):

lol your* sorry